Glenn and Paula want to dig an irrigation channel for their garden. Together it takes them 1.8 hours to dig an appropriately sized irrigation channel. If Paula can dig the channel by herself in 2.2 hours, how long would it take Glenn to dig the channel by himself? Round to the nearest tenth, if necessary.
Answer by mananth(16949)

together they take 1.8 hours= 1 8/10 = 1 4/5 = 9/5 hours
they do 5/9 of the job in 1 hour.
..
Paula alone does it 2.2 hours = 2 2/10 = 2 1/5 = 11/5 hours
she alone does 5/11 of the job in 1 hour.
..
5/9 -5/11 is the job done by Glenn alone in 1 hour
(55-45)/11
=10/11 of the job in 1 hour.
Glenn takes 11/10 hours to do the job alone = 1.1 hours